OpenStudy (anonymous):

A circle has an area of 441π cm2. What is the circumference

OpenStudy (anonymous):

area = pi r^2 441pi = pi r^2 r^2 = 441 r = sqrt(441) circumference= 2pi r circumfrence = 2 pi (sqrt 441) Hope this helps!

OpenStudy (anonymous):

\[A =\pi r^2\] \[C = 2\pi r\] \[441 = \pi r^2\]

OpenStudy (anonymous):

Solve for r, then use it to find C.

OpenStudy (anonymous):

Or nvm, I guess bob solved it for ya.

OpenStudy (anonymous):

a. 21π cm c. 42 cm b. 42π cm d. 21 cm

OpenStudy (anonymous):

You have everything you need, you can solve it now :)